OpenStudy (anonymous):

The temperature of a chemical reaction ranges between 40 degrees Celsius and 180 degrees Celsius. The temperature is at its lowest point when t = 0, and the reaction completes 1 cycle during a 12-hour period. What is a cosine function that models this reaction? f(t) = 70 cos 12t + 110 f(t) = 110 cos 12t + 70 f(t) = −70 cos pi over 6t + 110 f(t) = −110 cos pi over 6t + 70

OpenStudy (anonymous):

@Jacob902 @pooja195 @Hero

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Michele_Laino @dan815

OpenStudy (michele_laino):

the amplitude A of your function is: \[\Large A = \frac{{180 - 40}}{2} = ...?\]

OpenStudy (anonymous):

140/2= 70

OpenStudy (anonymous):

So A?

OpenStudy (michele_laino):

no, since the period of our function is 12 hours, so the shape og our function has to be this: \[\Large \cos \left( {\frac{{2\pi }}{{12}}t} \right)\]

OpenStudy (michele_laino):

of*

OpenStudy (anonymous):

It's really hard so I don't really know what to plug in for the shape of function

OpenStudy (michele_laino):

since at t=0, the value is 40, then we have this drawing: